# Quantum enhanced real-time sensing of protein binding process

### v1

@@ Abstract
Analyzing biokinetic processes play significant role in understanding fundamental cellular functions. Many physics based technologies used to study such processes are limited by the shot noise inherent to the coherent states of light. These technologies can greatly advantage by leveraging quantum probes to improve the sensitivity of measurements in cellular biology. Surface Plasmon Resonance (SPR) technique has been used effectively to achieve label-free, real-time measurements of protein binding kinetics that constitutes an important biological phenomenon occurring near the cell membrane. Here, we demonstrate the integration of this technique with the two-mode bright squeezed state having fewer fluctuations as compared to the coherent state to improve the sensitivity of measurement in studying a protein binding process. We show 4dB of squeezing as we record the signal-to-noise ratio as the function of time and is maintained throughout the binding process. The quantum advantage as shown in terms of squeezing is achieved despite the total absorption of $74\%$ from the source until the final detection after the sensor. Overall we provide the most practical setup for improving the sensitivity of the time-dependent measurements involved in various biological processes at the molecular level.
@@ I Introduction
Quantum sensing and metrology is a promising field proving its importance by improving the sensitivity of measurements in various fields ranging from phase shift measurements, force measurements, imaging, spectroscopy and more. It has been demonstrated that the current biosensing technologies can greatly advantage by leveraging quantum probes such as the entangled photon sources and squeezed states of light. It is possible to even reach the Heisenberg limit which is fundamental theoretical limit provided by quantum theory to enhance the spatial and temporal resolution in the field of bioimaging [1]. One specialized sub-area of quantum sensing combines the benefits of both the worlds, quantum mechanics and the plasmonic technology [2, 3, 4] useful in studying fundamental biochemical processes occurring at the cellular level. There exists many other technologies such as Super-resolution fluorescence microscopy [5]that involves labelling of the specific cellular components with fluorescent dyes, Total Internal Reflection–Fluorescence Correlation Spectroscopy (TIR-FCS) [6, 7] to observe events near the cell surface, nanoplasmonic biosensors[8, 9, 10] involving metallic nano-films or nano-structures (mainly gold), among others pushing the limits to improve the sensing capabilities for the molecular level study of the cellular processes.
@@
However, the Surface Plasmon Resonance (SPR) has proved to be a powerful technique in medical research mainly used to analyse events at or near a metal-dielectric interface, where the metal is typically gold film or naoparticles and the dielectric constitutes the analyte under study [11, 12, 13]. It has been used effectively to achieve label-free, real-time measurements of protein binding kinetics which constitutes an important dynamic process occurring near the cell membrane [14, 15, 16]. The SPR technique is based on sensing the minute changes in refractive index $\Delta n$ of the sample which in turn gives information about concentration, affinity and the kinetic parameters of the protein binding event. Understanding the dynamic processes such as protein binding is essential in learning how diseases develop at the molecular level and finding the potential treatment strategies [17]. Protein binding plays an important role in the area of drug design and delivery and studying this process requires more sophisticated tools to advance the field of medical research [18, 19, 20, 21].
@@
Most of the current bio-sensing techniques including SPR systems use coherent source of light for which the sensitivity of measuring the refractive index $n$ is such that $\Delta n\propto\frac{1}{\sqrt{N}}$, where N is the total number of photons interacting with the sample. This is called the Shot Noise Limit (SNL). It has been shown that the performance of SPR technique is highly dependent on the noise characteristics of the light and that the SPR sesnsors are reaching their theoretical limit [22, 23, 24]. One of the ways to improve the sensitivity of measurements is by increasing the number of photons. However this might not be always feasible due to the common issue of photo-damage where higher intensities of light might destroy or alter the sample under study[25]. An effective way to improve the signal-to-noise ratio (SNR) is by reducing the noise and this is where quantum probes prove extremely useful. Marino et al. demonstrated quantum enhancement of $56\%$ when compared with the corresponding classical configuration of state-of-the-art plasmonic sensor[26] using the squeezed light. Recently, it is extended to apply for an array of sensors simultaneously or measure multiple parameters independently[27]. A theoretical study by Mpofu et al. compares the advantage of using different quantum probes specifically for the Surface Plasmon Resonance (SPR) sensing technique [28]. A recent study by the same group utilizes the single photons generated via Spontaneous Parametric Down Conversion (SPDC) to probe the protein binding kinetics using the SPR sensor[29, 30]. Thus improving the SNR is the key to improving the sensitivity of current technology.
@@
In our work we consider the Two Mode Bright Squeezed State (TMBSS) because of its continuous wave nature allows for a higher number of photons as compared to the discrete quantum sources while still maintaining the quantum correlations between the two modes. It can enhance the sensing capabilities specifically by reducing the fluctuations associated with the intensity difference signal. Pooser et al. has demonstrated a quantum plasmonic sensor using the Krescthcmann configuration and the TMBSS source to measure the shift in SPR dip for samples with different refractive index [31]. However, no previous study has experimentally investigated the quantum advantage of TMBSS in the context of time dependent SPR sensing to measure the protein binding affinities, association rates, and dissociation rates.
@@
In this paper, we present the most practical setup to integrate the TMBSS along with the SPR sensor and give proof of principle experiment showing the quantum advantage which stays almost constant throughout the protein binding process. This setup can be used to measure the kinetic parameters with higher precision than its classical counterpart when working in the shot noise limited region. We first explain in detail the workings of the SPR sensor that detects minute changes in the refractive index of the sample in contact with the gold film. The sensitivity of measurement of refractive index is directly related to the sensitivity of measuring the kinetic parameters. We then explain the noise reduction achieved using the TMBSS in the form of squeezing. We describe the experimental setup along with the data processing performed to measure the SNR. We present our findings from plotting the SPR dip useful in deciding the incident angle on sensor to quantifying the SNR as a function of time as the BSA protein binds to the gold film on the sensor. Finally we comment on the challenges as well as the future scope for the setup described in this study.
@@ II Schematics of the quantum enhanced surface plasmon sensor
Surface Plasmon Resonance is a well-known technique in the field of bio sensing. It is based on the phenomenon where the surface plasmon (electron oscillations) at the metal-dielectric interface gets coupled with the evanescent wave generated due to an incident electromagnetic field [32]. This results in the formation of Transverse Magnetic (TM) wave confined to the metal-dielectric interface called as the Surface Plasmon Polariton (SPP) wave. Although the phenomenon could be observed using different coupling arrangements, the most common setup used is the Krescthcmann geometry as shown in Fig 1a. Here the prism, the gold film, and the dielectric medium form the apparatus useful in sensing the refractive index of the dielectric. The SPP wave propagates along the x-axis and decays exponentially on both sides of the interface along the z-axis as shown in Fig 1a.
@@
$k_{x}=\frac{\omega}{c}\sqrt{\frac{\epsilon_{2}(\omega)*\epsilon_{3}(\omega)}{\epsilon_{2}(\omega)+\epsilon_{3}(\omega)}}$ (1)
@@
$n_{1}\sin{\theta_{r}}=\sqrt{\frac{\epsilon_{2}(\omega)*\epsilon_{3}(\omega)}{\epsilon_{2}(\omega)+\epsilon_{3}(\omega)}}$ (3)
@@
Thus, at a particular angle of incidence ${\theta_{r}}$ after the critical angle the SPP are formed leading to the absorption of energy from the incident field. This is called the Surface Plasmon Resonance angle. As we plot the reflectivity of this sensor against the angle of incidence, we can see an absorption dip at the angle of resonance. Thus excitation of the SPP wave is highly sensitive to the refractive index of the dielectric medium at the interface. If we keep the thickness of the metal and the frequency of the laser constant, even the slightest change in the refractive index of the sample can be detected using the shift in the resonance dip. This is shown by the plot in Fig 1b which is based on the theoretical framework of electromagnetic wave propagating through a homogeneous film situated between two semi-finite media as explained in detail in the appendix section.
@@
This technique is already used to study molecular interactions and other parameters useful for bio-sensing. It can be used to measure the affinity of a particular sample, its concentration, specificity as well as the rate of interaction. Here we use this method in studying the rate at which the protein binds the gold film which serves as a preliminary proof of concept that could be extended to study other bio kinetic processes. If we lock the angle of incidence on the left side of the absorption dip, the changes in $n_{3}$ as the protein binds to the gold, could be detected by plotting the change in reflectivity as a function of time. This in turn could be translated to find the rate of the association and the dissociation of the protein as it binds and unbinds from the gold surface.
@@
Although an established technique, any protein binding parameter measured using a coherent source of light is shot noise limited. Even When all the technical noise is removed, the shot noise that is inherently related to the photon statistics limits the ability to distinguish two signals within a certain range limited by the fluctuations associated with the signal. The precision of the measurement of parameters studied in the protein binding analysis could be improved by using quantum probes such as the TMBSS source instead of the coherent source. Next we elaborate on the quantum advantage of using TMBSS light.
@@
The two-mode bright squeezed state referring to the intensity difference squeezing is a well-known quantum state that has fewer fluctuations as compared to the coherent state. It could be generated using different non linear process with atomic systems [33] being the most common. Here we use a Rubidium atomic vapour as a non-linear medium in which a coherent field and a vacuum field interact via the Four Wave Mixing (FWM) process to generate two quantum correlated beams namely the probe and the conjugate.
@@
Figure 2: a. Schematics of Surface Plasmon Resonance sensor using TMBSS light. The probe beam passing through the sensor undergoes an absorption due to SPP and the output Heisenberg operator is denoted by $a_{out}$. The conjugate beam intensity is adjusted according to the probe using an absorber and the associated Heisenberg operator is denoted by $b_{out}$. The resulting photo current is subtracted and constitutes the signal for the experiment given by $\langle{N}_{out}\rangle=\left\langle a_{out}^{\dagger}a_{out}\mathbin{-b_{out}^{\dagger}b_{out}}\right\rangle$. b. Schematics of SPR sensor using coherent light replicating the experimental conditions for probe and conjugate.
@@
As shown in Fig 2a. the two-mode bright squeezed state which ideally can be described as two single modes is defined using the Heisenberg picture given by the equations,
@@
where r is the squeezing parameter and the $a_{1},a_{in},a_{out}$ being the Heisenberg operators representing the electromagnetic field propagation as described in Fig 2a. The $r$ parameter in the context of the FWM process is related to the gain parameter $g$ such that,
$g=\cosh^{2}{r}$ (5)
where g is the ratio of the output probe intensity $\left\langle a_{1}^{\dagger}a_{1}\right\rangle$ and the input seed probe intensity $\left\langle a_{in}^{\dagger}a_{in}\right\rangle$. Notice that the input modes, the coherent state $\left|u\right\rangle$ and the vacuum state $\left|0\right\rangle$ are the single modes. Using the balanced detection technique as described in [34], the signal is given by $\langle{N}_{out}\rangle=\left\langle a_{out}^{\dagger}a_{out}\mathbin{-b_{out}^{\dagger}b_{out}}\right\rangle$ and the associated fluctuations are given by $\Delta N_{out}=\sqrt{\left\langle N_{out}^{2}\right\rangle-\left\langle N_{out}\right\rangle^{2}}$.
@@
To maintain the squeezing conditions before and after the sensor we match the absorption on the conjugate beam with that on the probe. Next we explain the experimental set up in detail used to study the protein binding process as an example of a biokientic process using the Two Mode Bright Squeezed State as well as using the coherent state.
@@ III Experimental set up
Figure 3: (a) Experimental setup showing the TMBSS generation using the Rubidium vapor and its application to study protein binding process using SPR sensor. (b) The $\chi^{(3)}$ FWM process inside a Rb85 atom (c) Schematic of the flow cell and the sensor arrangement.
The experimental setup we used for the quantum-enhanced protein binding analysis can be divided into three parts. Firstly we generate the two mode bright squeezed state using a non-linear process occurring inside a cell containing the Rb85 atomic vapour. This process generates two beams namely the probe and the conjugate that are quantum correlated. In the next step the probe beam passes through the Surface Plasmon Resonance (SPR) sensor where the light matter interaction at the sample-sensor interface enables the study of protein binding process. In the final step, the data is collected using the balanced detection method as theoretically described in the previous section. Each step is further explained below.

### v2

# Quantum enhanced real-time sensing of protein-gold adsorption kinetics

@@ Abstract
Analyzing the kinetics of biological processes plays a significant role in understanding fundamental cellular functions. Many physics-based technologies used to study such processes are limited by the shot noise inherent to the coherent states of light. These technologies can greatly benefit by leveraging quantum probes to improve the sensitivity of measurements in cellular biology. Surface Plasmon Resonance (SPR) technique has been used effectively to achieve label-free, real-time measurements of protein binding kinetics, which constitutes an important biological phenomenon occurring near the cell membrane. Here, we demonstrate the integration of this technique with the two-mode bright squeezed state having fewer fluctuations as compared to the coherent state to improve the sensitivity of measurement in studying a protein-gold adsorption process. We show 4dB of squeezing as we record the signal-to-noise ratio as the function of time and it is maintained throughout the kinetic process. The quantum advantage as shown in terms of squeezing is achieved despite the total absorption of $74\%$ from the source until the final detection after the sensor. Overall, we provide the most practical setup for improving the sensitivity of the time-dependent measurements involved in various biological processes at the molecular level.
@@ I Introduction
Quantum sensing and metrology is a promising field proving its importance by improving the sensitivity of measurements in various fields ranging from phase-shift measurements, force measurements, imaging, spectroscopy, and more. It has been demonstrated that the current biosensing technologies can greatly advantage by leveraging quantum probes such as the entangled photon sources and squeezed states of light. It is possible to even reach the Heisenberg limit which is fundamental theoretical limit provided by quantum theory to enhance the spatial and temporal resolution in the field of bioimaging [1]. One specialized sub-area of quantum sensing combines the benefits of both the worlds, quantum mechanics and the plasmonic technology [2, 3, 4] useful in studying fundamental biochemical processes occurring at the cellular level. There exists many other technologies such as super-resolution fluorescence microscopy [5]that involves labeling of the specific cellular components with fluorescent dyes, Total Internal Reflection–Fluorescence Correlation Spectroscopy (TIR-FCS) [6, 7] to observe events near the cell surface, nanoplasmonic biosensors[8, 9, 10] involving metallic nano-films or nano-structures (mainly gold), among others pushing the limits to improve the sensing capabilities for the molecular level study of the cellular processes.
@@
However, the Surface Plasmon Resonance (SPR) has proved to be a powerful technique in medical research mainly used to analyze events at or near a metal-dielectric interface, where the metal is typically gold film or nanoparticlesand the dielectric constitutes the analyte under study [11, 12, 13]. It has been used effectively to achieve label-free, real-time measurements of protein binding kinetics which constitutes an important dynamic process occurring near the cell membrane [14, 15, 16]. The SPR technique is based on sensing the minute changes in refractive index $\Delta n$ of the sample which in turn gives information about concentration, affinity and the kinetic parameters of the protein binding event. Understanding the dynamic processes such as protein binding is essential in learning how diseases develop at the molecular level and finding the potential treatment strategies [17]. Protein binding plays an important role in the area of drug design and delivery and studying this process requires more sophisticated tools to advance the field of medical research [18, 19, 20, 21].
@@
Most of the current bio-sensing techniques including SPR systems use coherent source of light for which the sensitivity of measuring the refractive index $n$ is such that $\Delta n\propto\frac{1}{\sqrt{N}}$, where N is the total number of photons interacting with the sample. This is called the Shot Noise Limit (SNL). It has been shown that the performance of SPR technique is highly dependent on the noise characteristics of the light and that the SPR sensors are reaching their theoretical limit [22, 23, 24]. One of the ways to improve the sensitivity of measurements is by increasing the number of photons. However, this might not be always feasible due to the common issue of photo-damage where higher intensities of light might destroy or alter the sample under study[25]. An effective way to improve the signal-to-noise ratio (SNR) is by reducing the noise and this is where quantum probes prove extremely useful. Marino et al. demonstrated quantum enhancement of $56\%$ when compared with the corresponding classical configuration of state-of-the-art plasmonic sensor[26] using the squeezed light. Recently, it is extended to apply for an array of sensors simultaneously or measure multiple parameters independently[27]. A theoretical study by Mpofu et al. compares the advantage of using different quantum probes specifically for the Surface Plasmon Resonance (SPR) sensing technique [28]. A recent study by the same group utilizes the single photons generated via Spontaneous Parametric Down Conversion (SPDC) to probe the protein-gold binding kinetics using the SPR sensor[29, 30]. Thus improving the SNR is the key to improving the sensitivity of current technology.
@@
In our work we consider the Two Mode Bright Squeezed State (TMBSS) because of its continuous wave nature allows for a higher number of photons as compared to the discrete quantum sources while still maintaining the quantum correlations between the two modes. It can enhance the sensing capabilities specifically by reducing the fluctuations associated with the intensity difference signal. Pooser et al. has demonstrated a quantum plasmonic sensor using the Krescthcmann configuration and the TMBSS source to measure the shift in SPR dip for samples with different refractive index [31]. However, no previous study has experimentally investigated the quantum advantage of TMBSS in the context of time dependent SPR sensing to measure the protein-gold adsorption kinetics. Protein adsorption on gold induces a response similar to the protein binding process as it changes the refractive index of the sample when the protein associates with or dissociates from the gold film. Any further references to protein binding to the gold film are written in the context of the adsorption process.
@@
In this paper, we present the most practical setup to integrate the TMBSS along with the SPR sensor and give proof of principle experiment showing the quantum advantage which stays almost constant throughout the protein adsorption process.Here, we mainly focus on the association of the protein with the gold surface. This setup can be used to measure the kinetic parameters with higher sensitivity than its classical counterpart when working in the shot noise limited region. We first explain in detail the workings of the SPR sensor that detects minute changes in the refractive index of the sample in contact with the gold film. The sensitivity of measurement of refractive index is directly related to the sensitivity of measuring the kinetic parameters. We then explain the noise reduction achieved using the TMBSS in the form of squeezing. We describe the experimental setup along with the data processing performed to measure the SNR. We present our findings from plotting the SPR dip useful in deciding the incident angle on sensor to quantifying the SNR as a function of time as the BSA protein binds to the gold film on the sensor. We show improvement in the sensitivity of estimating the observable rate constant of the binding process. Finally we discuss the challenges, the significance as well as the future implications of the setup described in this study.
@@ II Schematics of the quantum enhanced surface plasmon sensor
Surface Plasmon Resonance is a well-known technique in the field of bio sensing. It is based on the phenomenon where the surface plasmon (electron oscillations) at the metal-dielectric interface gets coupled with the evanescent wave generated due to an incident electromagnetic field [32, 33, 34]. This results in the formation of Transverse Magnetic (TM) wave confined to the metal-dielectric interface called as the Surface Plasmon Polariton (SPP) wave. Although the phenomenon could be observed using different coupling arrangements, the most common setup used is the Krescthcmann geometry as shown in Fig. 1(a). Here the prism, the gold film, and the dielectric medium form the apparatus useful in sensing the refractive index of the dielectric. The SPP wave propagates along the x-axis and decays exponentially on both sides of the interface along the z-axis as shown in Fig. 1(a).
@@
$k_{x}=\frac{\omega}{c}\sqrt{\frac{\epsilon_{2}(\omega)\cdot\epsilon_{3}(\omega)}{\epsilon_{2}(\omega)+\epsilon_{3}(\omega)}}$ (1)
@@
$n_{1}\sin{\theta_{r}}=\sqrt{\frac{\epsilon_{2}(\omega)\cdot\epsilon_{3}(\omega)}{\epsilon_{2}(\omega)+\epsilon_{3}(\omega)}}$ (3)
@@
Thus, at a particular angle of incidence ${\theta_{r}}$ after the critical angle the SPP are formed leading to the absorption of energy from the incident field. This is called the Surface Plasmon Resonance angle. As we plot the reflectivity of this sensor against the angle of incidence, we can see an absorption dip at the angle of resonance. Thus excitation of the SPP wave is highly sensitive to the refractive index of the dielectric medium at the interface. If we keep the thickness of the metal and the frequency of the laser constant, even the slightest change in the refractive index of the sample can be detected using the shift in the resonance dip. This is shown by the plot in Fig. 1(b) which is based on the theoretical framework called Transfer-matrix method [35] which describes propagation of an electromagnetic wave through a homogeneous film situated between two semi-infinite media as explained in detail in the appendix section. Note that the reflectance curve is more linear on the left side of the resonance angle as compared to the right side. Operating in this region produces higher shift in intensity measurement for a small shift in the resonance angle and avoids saturation of the measurements. As shown by the marker in Fig. 1(b) we operate on the left side of the resonance angle as it allows for a higher dynamic range in measurements accommodating various analyte concentrations without requiring frequent recalibration. This region offers more stability considering the mechanical vibrations or thermal drift allowing for more reproducible and reliable measurements.
@@
This technique is already used to study molecular interactions and other parameters useful for bio-sensing. It can be used to measure the affinity of a particular sample, its concentration, specificity, and the rate of interaction. Here we use this method to study the rate at which the protein adsorbs on to the gold film, which serves as a preliminary proof of concept that could be extended to study other biokinetic processes. If we lock the angle of incidence on the left side of the absorption dip, the changes in $n_{3}$ as the protein binds to the gold, could be detected by plotting the change in reflectivity as a function of time. This in turn could be translated to find the rate of the association and the dissociation of the protein as it binds and unbinds from the gold surface.
@@
Although an established technique, any protein binding parameter measured using a coherent source of light is shot noise limited. Even when all the technical noise is removed, the shot noise that is inherently related to the photon statistics limits the ability to distinguish two signals within a certain range limited by the fluctuations associated with the signal. The sensitivity of the estimation of parameters studied in the protein binding analysis could be improved by using quantum probes such as the TMBSS source instead of the coherent source. Next, we elaborate on the quantum advantage of using TMBSS light.
@@
The two-mode bright squeezed state referring to the intensity difference squeezing is a well-known quantum state that has fewer fluctuations compared to the coherent state. It could be generated using different non-linear processes with atomic systems [36] being the most common. Here we use a rubidium atomic vapor as a nonlinear medium in which a coherent field and a vacuum field interact via the Four-Wave Mixing (FWM) process to generate two quantum correlated beams namely the probe and the conjugate.
@@
Figure 2: (a) Schematics of Surface Plasmon Resonance sensor using TMBSS light. The probe beam passing through the sensor undergoes an absorption due to SPP and the output Heisenberg operator is denoted by $a_{out}$. The conjugate beam intensity is adjusted according to the probe using an absorber and the associated Heisenberg operator is denoted by $b_{out}$. The resulting photo current is subtracted and constitutes the signal for the experiment given by $\langle{N}_{out}\rangle=\left\langle a_{out}^{\dagger}a_{out}\mathbin{-b_{out}^{\dagger}b_{out}}\right\rangle$. (b) Schematics of SPR sensor using coherent light replicating the experimental conditions for probe and conjugate.
@@
As shown in Fig. 2(a), the two-mode bright squeezed state which ideally can be described as two single modes is defined using the Heisenberg picture given by the equations,
@@
where r is the squeezing parameter and the $a_{1},a_{in}$ being the Heisenberg operators representing the electromagnetic field propagation as described in Fig. 2(a). Notice that the input modes, the coherent state $\left|u\right\rangle$ and the vacuum state $\left|0\right\rangle$ are the single modes. Using the balanced detection technique as described in [37], the signal is given by $\langle{N}_{out}\rangle=\left\langle a_{out}^{\dagger}a_{out}\mathbin{-b_{out}^{\dagger}b_{out}}\right\rangle$ and the associated fluctuations are given by $\Delta N_{out}=\sqrt{\left\langle N_{out}^{2}\right\rangle-\left\langle N_{out}\right\rangle^{2}}$.
@@
To maintain the squeezing conditions before and after the sensor we match the absorption on the conjugate beam with that on the probe. Next we explain the experimental set up in detail used to study the protein-gold adsorption process as an example of a biokinetic process using the Two Mode Bright Squeezed State as well as using the coherent state.
@@ III Experimental set up
Figure 3: (a) Experimental setup showing the TMBSS generation using the Rubidium vapor and its application to study protein-gold binding process using SPR sensor. The solid lines after the Rb cell indicate the probe (orange) and the conjugate (green)beams. The intensity and the noise measurements at detector A and detector B are subtracted during the post-processing of the data. The flip mirrors indicated in gray are flipped up (active) only when performing the classical measurement. The classical beam (dotted line) splits into two beams at the Polarizing Beam Splitter (PBS) and the intensity is adjusted using the Half Wave Plate (HWP) to replicate the intensities in TMBSS case.
(b) The $\chi^{(3)}$ FWM process inside a Rb85 atom (c) Schematic of the flow cell and the sensor arrangement.
The experimental setup we used for the quantum-enhanced protein-gold binding analysis can be divided into three parts. Firstly we generate the two mode bright squeezed state using a non-linear process occurring inside a cell containing the Rb85 atomic vapor. This process generates two beams namely the probe and the conjugate that are quantum correlated. In the next step the probe beam passes through the Surface Plasmon Resonance (SPR) sensor where the light matter interaction at the sample-sensor interface enables the study of protein binding process. In the final step, the data is collected using the balanced detection method as theoretically described in the previous section. Each step is further explained below.